If the probabilities that a computer technician will repair 3,4,5,6,7, or 8 or more laptops on any given workday are 0.12, 0.19, 0.28, 0.24, 0.10, and 0.07, respectively, what is the probability that he will service at least 5 laptops on his next day at work?

Answer by ikleyn(52781)

.

P(at least 5) = P(5) + P(6) + P(7) + P(8 or more) = 0.28 + 0.24 + 0.10 + 0.07.